# Gradient correlation is needed to accelerate SGD with momentum

### v1

@@ F.1 Convex-Expectation
F.2 Strongly convex - Expectation
@@
F.3 Convex - almost sure
@@
Following the line of Sebbouh et al. (2021), Theorem 9 is the key result of our proof.
@@
F.4 Strongly convex - almost sure
Similarly to Section F.3, we extend statement (7) of Theorem 3 to get a new, asymptotically better, almost sure convergence result.
@@ F.4 Strongly convex - almost sure
Appendix G Proofs of Section 4
@@ G.5 Bound on RACOGA
Appendix H RACOGA values in linear regression
@@ H.1 Two functions in two dimensions
H.2 RACOGA is high over most of the space
In Section H.1, we considered a $2$-dimensional example with $2$-functions. Increasing dimension and adding functions, the problem of characterizing RACOGA values becomes harder. In the case of independent data, it is possible to give a theoretical result considering the RACOGA values in expectation over the data.
@@ H.2 RACOGA is high over most of the space
Proposition 9 indicates that when having a large amount of data, we can expect the RACOGA values to be positive over a large part of the space. The proof of Proposition 9 is deferred in Appendix H.4.
@@
H.3 The curvature problem: first order algorithms are attracted by low RACOGA value areas
In Section H.2, we showed that in the case $d>N$ where $N$ is not too small, one can expect RACOGA values to be high over most of the space. This statement is reinforced in the presence of correlation inside data (Figure 9(b)).
However, if $d$ is high, even though we sample a large amount of points to evaluate RACOGA values, we could miss non positive RACOGA value areas if these areas are too small. On Figure 10, we see that these areas indeed exist. Moreover, although they are tiny with respect to the whole space (Section H.2), deterministic algorithms, namely GD (Algorithm 6) and NAG (Algorithm 7), dive into these areas and stayed trapped inside. Strikingly, SNAG behaves differently and it manages not to get stuck in the same area.

### v2

@@ G.1 Convex-Expectation
G.2 Strongly convex - Expectation
@@
G.3 Convex - almost sure
@@
Following the scheme of the proof of Theorem 3.1 in Sebbouh et al. (2021), the Theorem 10 is the key result of our proof.
@@
G.4 Strongly convex - almost sure
Similarly to Section G.3, we extend statement (8) of Theorem 3 to get a new, asymptotically better, almost sure convergence result.
@@ G.4 Strongly convex - almost sure
Appendix H Proofs of Section 4
@@ H.5 Bound on RACOGA
Appendix I RACOGA values in linear regression
@@ I.1 Two functions in two dimensions
I.2 RACOGA is high over most of the space
In Section I.1, we considered a $2$-dimensional example with $2$-functions. Increasing dimension and adding functions, the problem of characterizing RACOGA values becomes harder. In the case of independent data, it is possible to give a theoretical result considering the RACOGA values in expectation over the data.
@@ I.2 RACOGA is high over most of the space
Proposition 9 indicates that when having a large amount of data, we can expect the RACOGA values to be positive over a large part of the space. The proof of Proposition 9 is deferred in Appendix I.4.
@@
I.3 The curvature problem: first order algorithms are attracted by low RACOGA value areas
In Section I.2, we showed that in the case $d>N$ where $N$ is not too small, one can expect RACOGA values to be high over most of the space. This statement is reinforced in the presence of correlation inside data (Figure 12(b)).
However, if $d$ is high, even though we sample a large amount of points to evaluate RACOGA values, we could miss non positive RACOGA value areas if these areas are too small. On Figure 13, we see that these areas indeed exist. Moreover, although they are tiny with respect to the whole space (Section I.2), deterministic algorithms, namely GD (Algorithm 6) and NAG (Algorithm 7), dive into these areas and stayed trapped inside. Strikingly, SNAG behaves differently and it manages not to get stuck in the same area.